OpenStudy (anonymous):

A(t)=t^5-5t^4+20t^3-17 I am having trouble setting the derivative equal to zero. Please help!

OpenStudy (anonymous):

What did you get for the derivative?

OpenStudy (anonymous):

I've got \[A' = 5t^4-20t^3 + 60t^2 = 5t^2(t^2-4t + 12)\]

OpenStudy (anonymous):

THat's exactly what I've got

myininaya (myininaya):

a*b=0 implies either a=0 or b=0 or both=0

OpenStudy (anonymous):

Well I know that 5t^2=0 but I when I solve the parentheses for 0 (using quadratic formula) it comes out negative. I can't fit that into the rest of my problem!

OpenStudy (anonymous):

(I am solving for the intervals of increasing and decreasing in the function)

myininaya (myininaya):

you mean the number under the square root is negative?

OpenStudy (anonymous):

Yes

myininaya (myininaya):

so it is imaginary that means your only critical number is 0

OpenStudy (anonymous):

That just means that that part of the equation doesn't have any real roots.

OpenStudy (anonymous):

Okay. So if I am solving for the intervals, do I just divide my number in half?

myininaya (myininaya):

I like using test points before and after each critical number

OpenStudy (anonymous):

You're looking for intervals for what? Where it's increasing and decreasing?

OpenStudy (anonymous):

Yes, @polpak

myininaya (myininaya):

the derivative tells us the slope therefore tells us if is increasing or decreasing so choose a number before and after 0 to see what if it is increasing or decreasing

OpenStudy (anonymous):

So it's increasing from negative infinity to infinity, correct?

OpenStudy (anonymous):

The derivative is always positive except at 0. So it is increasing for \(x \in (-\infty,0) \bigcup (0,\infty)\)

OpenStudy (anonymous):

It is not increasing at 0.

myininaya (myininaya):

it has a horizontal tangent at x=0 thats what we found when we set A'=0

myininaya (myininaya):

A'=0 means we are trying to find all x that has slope 0

OpenStudy (anonymous):

Okay, thanks guys!